Joe is 15 years older than Sally. in 15 years Joe will be 3 times as old as sally is now. How old is Sally?: Joe is 15 years older than Sally. in 15 years Joe will be 3 times as old as sally is now. How old is Sally?
Answer by vleith(924)

Let j be joe's age and s be sally's age.
Given j = s + 15
j - 15 = s
Also given
j+15 = 3s
Substituting from above yields
j + 15 = 3(j -15)
j + 15 = 3j - 45
60 = 2j
30 = j
Joe is 30
sally is 30-15 = 15